OpenStudy (anonymous):

The number of fish (f) in a fish tank increases when the number of algae blooms (a) decreases. Write the correct equation for this scenario, and solve for the number of fish when there are 4 algae blooms. Algae Blooms Fish 10 10 25 4

Mehek (mehek14):

algae = x fish = y \(\dfrac{4 - 10}{25 - 10} = \dfrac{-6}{15}\) \(y = \dfrac{-6}{15}x + b\) \(10 = \dfrac{-6}{15} * 10 + b\) \(10 = -4 + b\) \(6 = b\) \(y = \dfrac{-6}{15} + 6\)

Mehek (mehek14):

*\(y = \dfrac{-6}{15}x + 6\)

OpenStudy (anonymous):

f = 1a; f = 4 f = .1a; f = .4 f = 100 over a; f = 25 f = 100 over a; f = 400 these are the ansers they gave me and im so confused